A tunning fork produces 4 beats per second with another tunning fork of frequency 256 Hz. The first one is now loaded with a little wax and the beat frequency is found to increase to 6 per second. What was the original frequency of the tunning fork. 

Grade:9

3 Answers

Chetan Mandayam Nayakar
11 years ago

loading with wax decreases frequency. Therefore, original frequency=252 Hz

Prakhar Srivastava
11 years ago

DEAR Vrushabh

As Beat frequency =f1-f2=4 (given)

so is f2=256 f1 could be 252 or 260

Now f1 loaded so its frequency will increase

f1-f2=6

here it''s asked about original frequency therefore f1=260.

thanks.


Prakhar Srivastava
11 years ago

original frequency will be 252 hz..!!!